# BS-tree: A gapped data-parallel B-tree

### v1

@@ 4.1. Deletions
To delete a key, we first locate its position $i$ in a leaf node, using the equality search algorithm (discussed in Section 3.3). Then, we copy the key value from position $i+1$ to position $i$ and propagate it backwards to previous gap positions in the node. If $i$ is the last position in the leaf, we set $v.keys[i]$=MAXKEY.
@@
As in previous work (Wu et al., 2021; Rao and Ross, 2000), we do not take action for nodes with fewer than 50% occupied slots, as we anticipate insertions to be more frequent than deletions, so node merges or key redistributions are not expected to pay-off. If the last entry is deleted from a node, the node is marked as empty and the corresponding separator entry at its parent is ‘deleted’ by copying the next key into it.
@@
1 find leaf $v$ and position $r$ by running lines 1-3 of Alg 3
2if $v.keys[r]\neq k$ then
@@
5
6if $r==N-1$ then $\triangleright$ last key in node
@@ 7       $bitmap\leftarrow bitmap\oplus$ 0x0001
8       $replicasOfKey\leftarrow\_tzcnt(bitmap)$ for $i\leftarrow 0$ to $replicasOfKey-1$ do $\triangleright$ propagate backwards
@@
10
11else  $\triangleright$ $r$ is not the last position in the leaf
@@ 13       $replicasOfKey\leftarrow\_lzcnt(bitmap)$
14       $nextValidKey\leftarrow v.keys[r+replicasOfKey+1]$ for $i\leftarrow 0$ to $replicasOfKey$ do $\triangleright$ propagate backwards
@@ 4.2. Insertions
Inserting a new key $k$ to the BS-tree entails searching for the leaf node and the position in it to place it. Search is conducted by applying $succ_{>k}$ operations starting from the root and following the corresponding pointers. When we reach the leaf $v$ where in $k$ should be inserted, we apply a $succ_{\geq k}$ operation which finds the proper slot in $v$ to insert $k$. Then, we verify whether the slot $i$ returned by $succ_{\geq k}$ is occupied by another key. This is done by a simple test. If $v.keys[i]=v.keys[i+1]$, then we are sure that position $i$ is free, so we place $k$ there and finish. For example, assume that we want to insert key 55 to the leaf node of Figure 3. We apply a $succ_{\geq k}$ operation to the leaf, which will give us position 5. Since the next position (6) has the same key, position 5 corresponds to a free slot (gap), hence, we have directly put the inserted key 55 there. On the other hand, if the key at position $i$ is different compared to the key at position $i+1$, this means that the position is occupied. In this case, we first find the first position $j$ after $i$, which is unused (i.e., a gap), and right-shift all keys (and the corresponding record pointers) from position $i$ to position $j-1$, to make space, so that key $k$ can be inserted at position $i$. If there is no free position after $i$, then we move one position to the left (left-shift) all keys and record ids from position $i$ until the first free position to the left of $i$. Figure 4 (bottom) shows an example of inserting key 52. As the slot where 52 should go is occupied by 53 and it is not a gap (the key following 53 is not equal to 53), we search for the next gap, which is the position next to 53, right-shift 53 there and make room for the new key 52. Algorithm 6 describes BS-tree insertion.
@@
1 compute leaf $v$ and position $r$ by running lines 1-3 of Alg 3
2if $v.slotuse<N$ then
@@ 9             $keysForShift\leftarrow\_lzcnt(bitmap\ll r)$
12                   shift right $keysForShift$ keys of node $v$
@@ 17                   $keysForShift\leftarrow\_tzcnt(bitmap\gg(N-r-1))-1$
18                   shift left $keysForShift$ keys of node $v$
@@ 5. Key Compression
BS-tree, as it has been discussed so far, stores the exact keys in its nodes. Previous work on key compression for B+-tree (Bohannon et al., 2001) uses fixed-size partial keys. One issue with partial keys is the overhead of decompression which may compromise performance. For BS-tree, we opt for the frame-of-reference (FOR) compression approach, which has minimal decompression overhead.
@@ 6. Implementation Details
Node size and structure. As in previous work (Schlegel et al., 2009; Levandoski et al., 2013; Shahvarani and Jacobsen, 2016; Yan et al., 2019; Zhang et al., 2016; Kwon et al., 2023; Leis et al., 2013; Binna et al., 2018; Zhang et al., 2018; Ding et al., 2020a; Wu et al., 2021; Zhang et al., 2024), we aim at indexing large keys, each being a 64-bit unsigned integer. To take full advantage of our SIMD $succ_{>}$ implementation, each node stores a maximum of $N=16$ entries; based on this, we allocate $16\times 64=1024$ bits for the keys of each node. Hence, the keys of each node (internal or leaf) fill two cache lines (each cache line can store 64 bytes). This means that for $succ_{>}$, we perform 2 SIMD instructions per node by loading the keys at 2 registers of 512 bits (8 keys at each register). $1024$ bits are also allocated for keys in the compressed CBS-tree nodes; a compressed key array may have 32 32-bit or 64 16-bit entries.
@@ 7. Concurrency control
A wide range of concurrency control techniques has been proposed for the B+-tree and other indices, including lock coupling (Bayer and Schkolnick, 1977; Graefe, 2011), right-sibling pointers (Lehman and Yao, 1981), fine-grained locking with lock coupling and logical removals (Bronson et al., 2010), Bw-tree’s lock-free mechanism (Levandoski et al., 2013; Wang et al., 2018) and Read-Optimized Write Exclusion (ROWEX) (Leis et al., 2016). For B-trees, Leis et al. (Leis et al., 2019, 2016) proposed an Optimistic Lock Coupling (OLC) technique, which is easy to implement and highly efficient. In OLC, when a thread wants to read or modify a node, it first acquires an optimistic read lock, allowing it to traverse the tree while maintaining a local copy of the node’s state. If the thread intends to perform an update, it checks whether the node has been modified by another thread since it was read. If not, it applies the changes and commits them atomically. In the event of a conflict (i.e., if another thread has modified the node), the thread rolls back its changes and retries the operation from the root of the tree.
Our current implementation of BS-tree employs the OLC mechanism (Leis et al., 2019, 2016) with a slight modification in how node splits are handled. In the original OLC, when a thread takes on a write task, it splits the first full node (inner or leaf) it encounters during traversal. After completing the split, the thread restarts its traversal until it finds a path where every node has at least one empty slot. This efficiently reorganizes node contents, ensuring that insertions triggering splits at multiple tree levels are managed without requiring a global lock. The use of sparse nodes in BS-tree allows for a more flexible procedure. Specifically, when a split occurs at a leaf node, the restart process is only necessary if the separator must be stored in the first node of the couple, which has no available free slot. This relaxation reduces unnecessary re-traversals.
@@ 8. Experiments
We experimentally compare BS-tree to alternative main-memory indices (learned and non-learned). As in previous work (Wongkham et al., 2022; Binna et al., 2018; Leis et al., 2013), we have built and compared indices for key data only; record ids or references are not stored, but the objective of each index is to locate the position(s) of the searched key(s). The implementation of all methods is in C++ and compiled with gcc (v13) using the flags -O3 and -march=native. The experiments were conducted on a system with an 11th Gen Intel® Core™ i7-11700K processor with 8 cores, running at 3.60 GHz, 128 GB of RAM, having AVX 512 support. The operating system used was Ubuntu 22.04. We extended the codebase of GRE (Wongkham et al., 2022) to include BS-tree.
@@ 8.1. Setup
Datasets. We ran our tests on standard benchmarking real datasets used in previous work (Marcus et al., 2020; Kipf et al., 2019; Wongkham et al., 2022); each on consists of unsigned 64-bit integer keys. In Amazon BOOKS (Marcus et al., 2020; Kipf et al., 2019), each key represents the popularity of a specific book. In FB (Marcus et al., 2020; Kipf et al., 2019; Sandt et al., 2019), each key is a Facebook user-id. OSM (Marcus et al., 2020; Kipf et al., 2019) contains unique integer-encoded locations from OpenStreetMap. GENOME (Wongkham et al., 2022; Rao et al., 2014) includes loci pairs from human chromosomes. PLANET (Wongkham et al., 2022; goo, 2017), a planet-wide collection of integer-encoded geographic locations compiled by OpenStreetMap. According to (Wongkham et al., 2022; Zhang et al., 2024), OSM, FB, GENOME, and PLANET are complex real-world datasets that can pose challenges for learned indices. In contrast, the key distribution of BOOKS is easy to learn. We did not conduct experiments using synthetic datasets with common distributions, as, according to (Marcus et al., 2020), it would be trivial for a learned index to model such distributions. Although we do not include experiments with non-integer keys, BS-tree can be used for floats/doubles, by changing the SIMD intrinsics in Snippet 2, and for strings, after being dictionary encoded (see (Ding et al., 2020b)).
@@
Competitors. We compare our proposed BS-tree and its compressed version, denoted by CBS-tree, with five updatable learned and non-learned indices, for which the code was publicly available by the authors. We did not compare to methods found inferior in previous comparative studies (Binna et al., 2018; Wongkham et al., 2022) and to those with proprietary or unavailable code (e.g., (Kwon et al., 2023; Zhang et al., 2024)).
@@
Non-learned Indices. STX library (stx, 2013) is a fully optimized C++ implementation of a main-memory B+-tree. We use the set-based implementation from STX, which does not store values in the leaf nodes. For its construction, we used its fast bulk-loading method. We used the default block size of STX (256 bytes), so each leaf node holds 32 keys ($32\times 8=256$ bytes). We used two versions of the STX tree: the first is the original code, referred to as B+-tree, while the second version creates 25% empty space at the end of each leaf node, denoted by Sparse B+-tree (for fairness, as our BS-tree also proactively introduces 25% of gaps in leaves).
@@
We also compare to SIMD-based implementations of HOT (hot, 2018; Binna et al., 2018) and ART (Leis et al., 2013; art, 2022). These tries do not support bulk-loading. However, we found that pre-sorting the data improves the construction time for both and results in more efficient structures. The HOT code release does not support range queries, so we implemented them ourselves. The ART code also lacks range query support, but we were unable to implement it due to the trie’s complex structure. HOT cannot handle keys greater than $12^{63}-$ 1, so we removed values exceeding this limit from certain datasets.
@@
Learned Indices. ALEX (ale, 2020; Ding et al., 2020a) and LIPP (lip, 2021; Wu et al., 2021) are the state-of-the-art updatable learned indices for key-value pairs (Wongkham et al., 2022). To use them, for each dataset we used as value of each key the key itself. ALEX and LIPP both support bulk-loading. Neither uses SIMD intrinsics during search. The reason is their large leaves (in the order of KB), where SIMD-based search is not effective (see Fig. 2). ALEX uses exponential search in leaves, while LIPP eliminates the last-mile search entirely by ensuring that predicted positions are exact for each key in the index.
@@ 8.2. Construction Cost and Memory Footprint
In conclusion, the BS-tree has low construction cost, with the B+-tree exhibiting comparable performance alongside a very small memory footprint. Additionally, the CBS-tree achieves the fastest construction time and consumes from 56% to 94% less memory than all methods in FB, GENOME, and PLANET. Our results align with the findings of Wongkham et al. (Wongkham et al., 2022), which conclude that memory efficiency is not a distinct advantage of updatable learned indices.
@@ 8.4. Performance Counters
Besides throughput, we also compared all methods with respect to various performance counters, including instructions executed, cycles, mispredicted branches, and misses in L1 and TLB (Translation Lookaside Buffer) misses. As representative datasets, we selected BOOKS and FB. [^5] Recall that ALEX presents competitive behavior on BOOKS, whereas on FB BS-tree and CBS-tree have a large performance gap to other methods. We chose to measure Workload C (reads - writes), as it is update-heavy and the unpredictable nature of writes can lead to numerous splits, stressing the indices. To calculate these metrics, we utilized Leis’s perf_event code (per, 2018). The average performance measures per operation for all methods are presented in Table 3 and Table 4 for BOOKS and FB, respectively.
@@ 8.6. Impact of BS-tree design
In the final set of experiments, we assess the impact of the design choices in BS-tree. Figure 13 shows the impact of our data-parallel implementation using gaps that copy the key values from the next used slot, on the cost of writes. We compare our BS-tree implementation to an implementation of BS-tree, where nodes have gaps simply supported by a bitmap (as in (Ding et al., 2020a)). In this implementation, branching in each node is done by linear scan using the bitmap to ignore unused slots. [^6] As nodes are small, linear scan is better than binary search. Besides, binary and exponential search require values in gaps that obey the total order (Ding et al., 2020a). The figure shows that our SIMD implementation with gaps reduces the cost of insertions (and node searches, which are part of the insertion procedure) by 30%-35%.
@@ 9. Conclusions
We proposed BS-tree, a main-memory B+-tree that uses a simple and intuitive data-parallel implementation for branching at each level during search and updates. BS-tree is based on a novel implementation of gaps in nodes by duplicating existing keys, which does not affect SIMD-based branchless search in nodes. We use FOR compression mechanism for nodes that include keys with small differences. Our experimental evaluation demonstrates the superiority of BS-tree compared to open-source state-of-the-art non-learned and learned indices, with respect to construction time, memory footprint, and throughput for various workloads that include queries and updates in single- and multi-threaded processing. Inspired by (Shahvarani and Jacobsen, 2016), in the future, we plan to implement BS-tree in a hybrid setup, where the top (and infrequently updated levels) are handled by the GPU that allows much higher data parallelism and the lower (frequently updated) levels are handled by the CPU. In addition, we will study the support of string keys in BS-tree (one solution is to use Binary, ASCII or Base64 encoding (Ding et al., 2020a)).

### v2

@@ 4.1. Deletions
To delete a key, we first locate its position $i$ in a leaf node, using the equality search algorithm (discussed in Section 3.3). Then, we copy the key value from position $i+1$ to position $i$ and propagate it backwards to previous gap positions in the node. If $i$ is the last position in the leaf, we set v.keys[i]v.keys[i]=MAXKEY.
@@
As in previous work (WuZCCWX21; RaoR00), we do not take action for nodes with fewer than 50% occupied slots, as we anticipate insertions to be more frequent than deletions, so node merges or key redistributions are not expected to pay-off. If the last entry is deleted from a node, the node is marked as empty and the corresponding separator entry at its parent is ‘deleted’ by copying the next key into it.
@@
1
21exfind leaf $v$ and position $r$ by running lines 1-3 of Alg 3
3if $v.keys[r]\neq k$ then
@@
7
8if r==N1r==N-1 then $\triangleright$ last key in node
@@ 9  $bitmap\leftarrow bitmap\oplus$ 0x0001
10  $replicasOfKey\leftarrow\_tzcnt(bitmap)$
11 1exfor $i\leftarrow 0$ to replicasOfKey1replicasOfKey-1 do $\triangleright$ propagate backwards
@@ 10  $replicasOfKey\leftarrow\_tzcnt(bitmap)$
13
14else $\triangleright$ $r$ is not the last position in the leaf
@@ 16  $replicasOfKey\leftarrow\_lzcnt(bitmap)$
17  $nextValidKey\leftarrow v.keys[r+replicasOfKey+1]$
18 1exfor $i\leftarrow 0$ to replicasOfKeyreplicasOfKey do $\triangleright$ propagate backwards
@@ 4.2. Insertions
Inserting a new key $k$ to the BS-tree entails searching for the leaf node and the position in it to place it. Search is conducted by applying $succ_{>k}$ operations starting from the root and following the corresponding pointers. When we reach the leaf $v$ where in $k$ should be inserted, we apply a $succ_{\geq k}$ operation which finds the proper slot in $v$ to insert $k$. Then, we verify whether the slot $i$ returned by $succ_{\geq k}$ is occupied by another key. This is done by a simple test. If $v.keys[i]=v.keys[i+1]$, then we are sure that position $i$ is free, so we place $k$ there and finish. For example, assume that we want to insert key 55 to the leaf node of Figure 3. We apply a $succ_{\geq k}$ operation to the leaf, which will give us position 5. Since the next position (6) has the same key, position 5 corresponds to a free slot (gap), hence, we have directly put the inserted key 55 there. On the other hand, if the key at position $i$ is different compared to the key at position $i+1$, this means that the position is occupied. In this case, we first find the first position $j$ after $i$, which is unused (i.e., a gap), and right-shift all keys (and the corresponding record pointers) from position $i$ to position j1j-1, to make space, so that key $k$ can be inserted at position $i$. If there is no free position after $i$, then we move one position to the left (left-shift) all keys and record ids from position $i$ until the first free position to the left of $i$. Figure 4 (bottom) shows an example of inserting key 52. As the slot where 52 should go is occupied by 53 and it is not a gap (the key following 53 is not equal to 53), we search for the next gap, which is the position next to 53, right-shift 53 there and make room for the new key 52. Algorithm 6 describes BS-tree insertion.
@@
1
21excompute leaf $v$ and position $r$ by running lines 1-3 of Alg 3
3if $v.slotuse<N$ then
@@ 10     $keysForShift\leftarrow\_lzcnt(bitmap\ll r)$
13        shift right keysForShiftkeysForShift keys of node $v$
@@ 18        $keysForShift\leftarrow\_tzcnt(bitmap\gg(N-r-1))-1$
19        shift left keysForShiftkeysForShift keys of node $v$
@@ 5. Key Compression
BS-tree, as it has been discussed so far, stores the exact keys in its nodes. Previous work on key compression for B+-tree (BohannonMR01) uses fixed-size partial keys. One issue with partial keys is the overhead of decompression which may compromise performance. For BS-tree, we opt for the frame-of-reference (FOR) compression approach, which has minimal decompression overhead.
@@ 6. Implementation Details
Node size and structure. As in previous work (SchlegelGL09; LevandoskiLS13a; ShahvaraniJ16; YanLPZ19; ZhangAPKMS16; KwonLNNPCM23; LeisK013; BinnaZPSL18; ZhangLLAKKP18; DingMYWDLZCGKLK20; WuZCCWX21; ZhangQYB24), we aim at indexing large keys, each being a 64-bit unsigned integer. To take full advantage of our SIMD $succ_{>}$ implementation, each node stores a maximum of $N=16$ entries; based on this, we allocate $16\times 64=1024$ bits for the keys of each node. Hence, the keys of each node (internal or leaf) fill two cache lines (each cache line can store 64 bytes). This means that for $succ_{>}$, we perform 2 SIMD instructions per node by loading the keys at 2 registers of 512 bits (8 keys at each register). $1024$ bits are also allocated for keys in the compressed CBS-tree nodes; a compressed key array may have 32 32-bit or 64 16-bit entries.
@@ 7. Concurrency control
A wide range of concurrency control techniques has been proposed for the B+-tree and other indices, including lock coupling (BayerS77; Graefe11), right-sibling pointers (LehmanY81), fine-grained locking with lock coupling and logical removals (BronsonCCO10), Bw-tree’s lock-free mechanism (LevandoskiLS13a; WangPLLZKA18) and Read-Optimized Write Exclusion (ROWEX) (LeisSK016). For B-trees, Leis et al. (LeisH019; LeisSK016) proposed an Optimistic Lock Coupling (OLC) technique, which is easy to implement and highly efficient. In OLC, when a thread wants to read or modify a node, it first acquires an optimistic read lock, allowing it to traverse the tree while maintaining a local copy of the node’s state. If the thread intends to perform an update, it checks whether the node has been modified by another thread since it was read. If not, it applies the changes and commits them atomically. In the event of a conflict (i.e., if another thread has modified the node), the thread rolls back its changes and retries the operation from the root of the tree.
Our current implementation of BS-tree employs the OLC mechanism (LeisH019; LeisSK016) with a slight modification in how node splits are handled. In the original OLC, when a thread takes on a write task, it splits the first full node (inner or leaf) it encounters during traversal. After completing the split, the thread restarts its traversal until it finds a path where every node has at least one empty slot. This efficiently reorganizes node contents, ensuring that insertions triggering splits at multiple tree levels are managed without requiring a global lock. The use of sparse nodes in BS-tree allows for a more flexible procedure. Specifically, when a split occurs at a leaf node, the restart process is only necessary if the separator must be stored in the first node of the couple, which has no available free slot. This relaxation reduces unnecessary re-traversals.
@@ 8. Experiments
We experimentally compare BS-tree to alternative main-memory indices (learned and non-learned). As in previous work (WongkhamLLZLW22; BinnaZPSL18; LeisK013), we have built and compared indices for key data only; record ids or references are not stored, but the objective of each index is to locate the position(s) of the searched key(s). The implementation of all methods is in C++ and compiled with gcc (v13) using the flags -O3 and -march=native. The experiments were conducted on a system with an 11th Gen Intel® Core™ i7-11700K processor with 8 cores, running at 3.60 GHz, 128 GB of RAM, having AVX 512 support. The operating system used was Ubuntu 22.04. We extended the codebase of GRE (WongkhamLLZLW22) to include BS-tree.
@@ 8.1. Setup
Datasets. We ran our tests on standard benchmarking real datasets used in previous work (MarcusKRSMK0K20; sosd-neurips; WongkhamLLZLW22); each on consists of unsigned 64-bit integer keys. In Amazon BOOKS (MarcusKRSMK0K20; sosd-neurips), each key represents the popularity of a specific book. In FB (MarcusKRSMK0K20; sosd-neurips; SandtCP19), each key is a Facebook user-id. OSM (MarcusKRSMK0K20; sosd-neurips) contains unique integer-encoded locations from OpenStreetMap. GENOME (WongkhamLLZLW22; rao20143d) includes loci pairs from human chromosomes. PLANET (WongkhamLLZLW22; goo2017), a planet-wide collection of integer-encoded geographic locations compiled by OpenStreetMap. According to (WongkhamLLZLW22; ZhangQYB24), OSM, FB, GENOME, and PLANET are complex real-world datasets that can pose challenges for learned indices. In contrast, the key distribution of BOOKS is easy to learn. We did not conduct experiments using synthetic datasets with common distributions, as, according to (MarcusKRSMK0K20), it would be trivial for a learned index to model such distributions. Although we do not include experiments with non-integer keys, BS-tree can be used for floats/doubles, by changing the SIMD intrinsics in Snippet 2, and for strings, after being dictionary encoded (see (DingNAK20)).
@@
Competitors. We compare our proposed BS-tree and its compressed version, denoted by CBS-tree, with five updatable learned and non-learned indices, for which the code was publicly available by the authors. We did not compare to methods found inferior in previous comparative studies (BinnaZPSL18; WongkhamLLZLW22) and to those with proprietary or unavailable code (e.g., (KwonLNNPCM23; ZhangQYB24)).
@@
Non-learned Indices. STX library (stx_code) is a fully optimized C++ implementation of a main-memory B+-tree. We use the set-based implementation from STX, which does not store values in the leaf nodes. For its construction, we used its fast bulk-loading method. We used the default block size of STX (256 bytes), so each leaf node holds 32 keys ($32\times 8=256$ bytes). We used two versions of the STX tree: the first is the original code, referred to as B+-tree, while the second version creates 25% empty space at the end of each leaf node, denoted by Sparse B+-tree (for fairness, as our BS-tree also proactively introduces 25% of gaps in leaves).
@@
We also compare to SIMD-based implementations of HOT (hot_code; BinnaZPSL18) and ART (LeisK013; art_code). These tries do not support bulk-loading. However, we found that pre-sorting the data improves the construction time for both and results in more efficient structures. The HOT code release does not support range queries, so we implemented them ourselves. The ART code also lacks range query support, but we were unable to implement it due to the trie’s complex structure. HOT cannot handle keys greater than $2^{63}-1$, so we removed values exceeding this limit from certain datasets.
@@
Learned Indices. ALEX (alex_code; DingMYWDLZCGKLK20) and LIPP (lipp_code; WuZCCWX21) are the state-of-the-art updatable learned indices for key-value pairs (WongkhamLLZLW22). To use them, for each dataset we used as value of each key the key itself. ALEX and LIPP both support bulk-loading. Neither uses SIMD intrinsics during search. The reason is their large leaves (in the order of KB), where SIMD-based search is not effective (see Fig. 2). ALEX uses exponential search in leaves, while LIPP eliminates the last-mile search entirely by ensuring that predicted positions are exact for each key in the index.
@@ 8.2. Construction Cost and Memory Footprint
In conclusion, the BS-tree has low construction cost, with the B+-tree exhibiting comparable performance alongside a very small memory footprint. Additionally, the CBS-tree achieves the fastest construction time and consumes from 56% to 94% less memory than all methods in FB, GENOME, and PLANET. Our results align with the findings of Wongkham et al. (WongkhamLLZLW22), which conclude that memory efficiency is not a distinct advantage of updatable learned indices.
@@ 8.4. Performance Counters
Besides throughput, we also compared all methods with respect to various performance counters, including instructions executed, cycles, mispredicted branches, and misses in L1 and TLB (Translation Lookaside Buffer) misses. As representative datasets, we selected BOOKS and FB. [^5] Recall that ALEX presents competitive behavior on BOOKS, whereas on FB BS-tree and CBS-tree have a large performance gap to other methods. We chose to measure Workload C (reads - writes), as it is update-heavy and the unpredictable nature of writes can lead to numerous splits, stressing the indices. To calculate these metrics, we utilized Leis’s perf_event code (perfevent_code). The average performance measures per operation for all methods are presented in Table 3 and Table 4 for BOOKS and FB, respectively.
@@ 8.6. Impact of BS-tree design
In the final set of experiments, we assess the impact of the design choices in BS-tree. Figure 13 shows the impact of our data-parallel implementation using gaps that copy the key values from the next used slot, on the cost of writes. We compare our BS-tree implementation to an implementation of BS-tree, where nodes have gaps simply supported by a bitmap (as in (DingMYWDLZCGKLK20)). In this implementation, branching in each node is done by linear scan using the bitmap to ignore unused slots. [^6] As nodes are small, linear scan is better than binary search. Besides, binary and exponential search require values in gaps that obey the total order (DingMYWDLZCGKLK20). The figure shows that our SIMD implementation with gaps reduces the cost of insertions (and node searches, which are part of the insertion procedure) by 30%-35%.
@@ 9. Conclusions
We proposed BS-tree, a main-memory B+-tree that uses a simple and intuitive data-parallel implementation for branching at each level during search and updates. BS-tree is based on a novel implementation of gaps in nodes by duplicating existing keys, which does not affect SIMD-based branchless search in nodes. We use FOR compression mechanism for nodes that include keys with small differences. Our experimental evaluation demonstrates the superiority of BS-tree compared to open-source state-of-the-art non-learned and learned indices, with respect to construction time, memory footprint, and throughput for various workloads that include queries and updates in single- and multi-threaded processing. Inspired by (ShahvaraniJ16), in the future, we plan to implement BS-tree in a hybrid setup, where the top (and infrequently updated levels) are handled by the GPU that allows much higher data parallelism and the lower (frequently updated) levels are handled by the CPU. In addition, we will study the support of string keys in BS-tree (one solution is to use Binary, ASCII or Base64 encoding (DingMYWDLZCGKLK20)).